OpenStudy (anonymous):

write an equation in the form y=mx+b for the line that passes through (-6,9) and has a slope 1/2

pooja195 (pooja195):

\[\huge~y-y_1=m(x-x_1)\]

pooja195 (pooja195):

(x1,y1) m=slope

OpenStudy (anonymous):

2,9

OpenStudy (anonymous):

and soory for the hard time have much to learn

pooja195 (pooja195):

Its okay :) \[\huge~y-9=1/2(x+6)\]

pooja195 (pooja195):

Distribute \[1/2(x+6) \] 1/2 * x = 1/2x 1/2 * 6=?

OpenStudy (anonymous):

dont mind the top the answer is 3

pooja195 (pooja195):

good now we have y−9=1/2x+3 3+9=?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

3+9=12

pooja195 (pooja195):

good so we get y=1/2x+12

OpenStudy (anonymous):

thank you for everything that why im your fan

pooja195 (pooja195):

aww thanks :) and you are welcome! D